Resistance

Current

Voltage

𝑅1

3kΞ©

πŸ“ 𝒙 πŸπŸŽβˆ’πŸ’π‘¨

1.5 V

𝑅2

10 kΞ©

πŸ“ 𝒙 πŸπŸŽβˆ’πŸ’π‘¨

5V

𝑅3

5kΞ©

πŸ“ 𝒙 πŸπŸŽβˆ’πŸ’π‘¨

2.5 V

Total

18 kΞ©

πŸ“ 𝒙 πŸπŸŽβˆ’πŸ’π‘¨


1
Expert's answer
2021-03-18T14:04:43-0400

Given,

"R_1=3k\Omega"

"I_1=5\times 10^{-4}A"

"V_1=1.5V"


"R_2=10k\Omega"

"I_2=5\times 10^{-4}A"

"V_2=5V"


"R_3=5k\Omega"

"I_3=5\times 10^{-4}A"

"V_3=2.5V"


As here we can see that the current in all the three resistance is same and the potential at all the three resistance is different, hence we can conclude that all the three resistances are connected in series.

Hence, equivalent resistance "(R_{eq})=R_1+R_2+R_3"

"=3 k\Omega+ 10k\Omega+ 5k\Omega"

"=18k\Omega"

Net applied potential "N_{net}=V_1+V_2+V_3"

"=1.5V+5V+2.5V"

"=9V"